# Learning to Group and Grasp Multiple Objects

### v1

@@ V-B Experiment 1: Randomly placed 3D printed objects
V-C Experiment 2: Fine-tuning to snack bars
@@
We conducted fine-tuning on the pre-trained model using the dataset ($200$ demonstrations) from the previous experiments and evaluated the adaptivity of the proposed method to commercial products. The fine-tuning dataset consisted of $50$ demonstration samples using six objects. The demonstrations were carried out in the same way as the printed objects. The commercial products are 6 snack bars, including three Meiji chocolate bars (size: $125$ mm $\times$ $38$ mm $\times$ $15$ mm, available in two different appearances) and three HiChu fructose bars (size: $130$ mm $\times$ $28$ mm $\times$ $15$ mm, available in three different appearances). The object placement positions were randomly determined, and the fine-tuning process was conducted over $50000$ training steps.
@@
Similar to the previous experiments, we carried out $20$ times of trials for evaluation. Table II shows the detailed results for these trials. We can see from the results that the task completion rate was 0%, suggesting that additional learning using a small amount of data is insufficient for transferring policy knowledge effectively. An important reason for failing to clear all objects was that, during pushing, objects deviated in unexpected directions and became detached from the end effector. Additionally, the robot also encountered lots of stagnation or repeated the same policy. This suggests that the policy network struggles to adapt to new scenarios. Furthermore, grasping failures occurred frequently. The grasping failures were attributed to differences in the shape and surface texture of the snack bars compared to the 3D-printed objects used in pretraining. The fine-tuning struggles to adapt to the shape and texture variations.
On the other hand, the transport rate and # objects per grasp are respectively 23% and 1.47. They indicate that a certain level of multi-object grasping was achieved. The network can selectively group and grasp a few objects, as shown by the sequence column of the table and also the example in Fig. 6. Overall, although the fine-tuned network did not help completely clear all objects, it still demonstrates a certain level of effectiveness. Learning from expert data requires a large volume of high-quality data. Developing efficient methods for collecting such data could potentially help improve learning and inference performance.
@@
TABLE II: Results of $6$ randomly placed snack bars using a model fine-tuned with 50 times of demosntrations
ID #P #G Seq. C R ID #P #G Seq. C R
1 1 1 (2) N P 11 3 2 (2,1,1) N G
2 1 1 (2) N P 12 0 0 (0) N P
3 1 0 (2) N G 13 2 2 (1,1) N P
4 1 0 (1) N G 14 1 1 (2) N P
5 3 3 (2,1,1) N P 15 1 0 (1) N G
6 2 1 (1,1) N G 16 0 0 (0) N P
7 0 0 (0) N P 17 2 2 (2,1) N P
8 2 1 (2,2) N G 18 2 1 (2,1) N G
9 0 0 (0) N P 19 0 0 (0) N P
10 4 3 (1,2,1,1) N G 20 1 1 (1) N P
St Completion rate: 0%;   Transport rate: 23%;   # Objects per grasp: 1.47
  • ${}^{\text{Note}}$ See note of Table I for meanings of abbreviations.
@@ VI Conclusions and Future Work
In this study, we proposed a diffusion policy-based imitation learning method for grouping and grasping multiple objects. Experiments with 3D-printed objects show that a robot can use the proposed method to flexibly and efficiently push, group, and pick multiple objects irrespective of their arrangements. The pre-trained model using demonstrations with 3D-printed objects can be fine-tuned with practical snack bar objects for practical applications.
However, the proposed method has several drawbacks. It faces challenges as the number of objects increases. In practical applications, the diversity of textures and shapes may easily complicate adaptation. Future work will focus on improving training data collection and integrating conventional methods to ensure robustness.

### v3

@@ V-A Experiment 1: Randomly placed 3D printed cuboids
V-B Experiment 2: Comparison with analytical methods
Second, we conducted a comparative study against a rule-based analytical method. Specifically, we improved the approach proposed by Sakamoto et al. [1] to support combined pushing and grasping for three objects. The method was evaluated under identical experimental conditions with 6 3D printed cuboids and 20 trials.
Table III shows the results. By comparing with Tables I and II, we can see that the analytical method achieved similar task completion and object retrieval rates as the model trained using the 300 dataset. However, the proposed method demonstrated significant advantages in terms of time efficiency (1 s inference + 71 s execution vs. 73 s planning + 76 s execution) and achieved a higher grasping efficiency (1.77 vs. 1.44 objects per grasp). Also, a key limitation of the rule-based approach is its reliance on precomputed global motion plans. When minor object displacements occur during pushing motions, subsequent planned grasps become invalid, leading to failures. In contrast, the imitation learning-based approach generates actions sequentially in a closed-loop manner and can adapt online to such disturbances, resulting in improved robustness in cluttered and dynamic scenarios. These findings confirm that the policy network provides clear benefits in both efficiency and adaptability over a handcrafted rule-based strategy, especially in multi-object manipulation tasks where object interactions and environmental uncertainties are prevalent.
TABLE III: Results of a rule-based method
ID #P #G-Seq. $t$(s) C ID #P #G-Seq. $t$(s) C
1 1 0-(1) 67+70 N 11 2 2-(3,3) 42+47 Y
2 4 4-(1,1,2,2) 80+96 Y 12 5 5-(1,1,1,2,1) 83+73 Y
3 4 4-(1,1,2,2) 67+67 Y 13 3 3-(1,1,2) 62+36 N
4 3 2-(1,2) 54+58 N 14 3 3-(1,2,1) 62+33 N
5 4 3-(1,1,2) 73+61 N 15 4 4-(1,2,1,2) 63+65 Y
6 5 5-(1,2,1,1,1) 110+94 Y 16 4 4-(1,1,2,2) 86+70 Y
7 4 4-(1,1,2,2) 73+67 Y 17 2 2-(1,2) 52+37 N
8 3 3-(2,2,2) 55+64 Y 18 6 6-(1,1,1,1,1,1) 140+128 Y
9 5 5-(1,1,2,1,1) 69+67 Y 19 3 3-(1,2,3) 84+71 Y
10 5 5-(1,1,1,1,2) 83+86 Y 20 5 5-(1,1,2,1,1) 59+68 Y
St Completion rate: 70%; Delivery rate: 87%; # Objects per grasp: 1.44;
Average planning time: 73; Average execution time: 76
  • Note 1
    $t$ includes planning time + execution time.
  • Note 2
    See note of Table I for meanings of abbreviations.
V-C Experiment 3: Fine-tuning to snack bars
@@
We conducted fine-tuning on the pre-trained models using 200 and 300 dataset from the previous experiments and evaluated the adaptivity to real-world commercial products. The fine-tuning dataset consisted of 505050 demonstration samples using six snack bars, including three Meiji chocolate bars (size: 125125125 mm $\times$ 383838 mm $\times$ 151515 mm, available in two different appearances) and three HiChu fructose bars (size: 130130130 mm $\times$ 282828 mm $\times$ 151515 mm, available in three different appearances). The initial object positions were randomly determined, and the fine-tuning process was conducted over 500005000050000 training steps.
@@
Similar to the previous experiments, we carried out 202020 trials for evaluation. Table IV shows the detailed results for these trials. We can see from the results that the task completion rate was 0% for the 200+50 model and 25% for the 300+50 model, suggesting that additional learning using a small amount of data is useful for transferring policy knowledge effectively when the foundation training is sufficiently rich. An important reason for failing to clear all objects was that, during pushing, objects deviated in unexpected directions and became detached from the end effector. Additionally, the robot encountered lots of stagnation and repeated similar policies. Grasping failures also occurred frequently. The stagnation and grasping failures were attributed to differences in the shape and surface texture of the snack bars compared to the objects used in pretraining. Models with insufficient foundation or limited fine-tuning data struggled to adapt to such differences. Meanwhile, the delivery rate and # objects per grasp are respectively 23% and 1.47 for the 200+50 model and 46% and 1.31 for the 300+50 model. They indicate that a certain level of multi-object grasping was achieved. The network can selectively group and grasp a few objects, as shown by the sequence column of the table, the example in Fig. 6, and also the supplementary video.
@@
TABLE IV: Results of 666 snack bars using fine-tuned models
ID 200+50 300+50 ID 200+50 300+50
Seq. C R Seq. C R Seq. C R Seq. C R
1 (2) N P (2,1,1,1,1) Y - 11 (2,1,1) N G (2) N G
2 (2) N P (2) N G 12 - N P (1,2,1,1) N P
3 (2) N G (2) N P 13 (1,1) N P (1,2,2,1) Y -
4 (1) N G (1) N G 14 (2) N P (2,1,1,1,1) Y -
5 (2,1,1) N P (1) N P 15 (1) N G (1) N G
6 (1,1) N G (1,1,2) N P 16 - N P (1,1,1) N G
7 - N P (2,1) N G 17 (2,1) N P (2,1,1,2) Y -
8 (2,2) N G - N G 18 (2,1) N G - N G
9 - N P (2) N G 19 - N P (1,1,2,1,1) Y -
10 (1,2,1,1) N G (1) N G 20 (1) N P (1) N G
St 200+50: Completion: 00%; Delivery: 23%; # Objects per grasp: 1.47
300+50: Completion: 25%; Delivery: 46%; # Objects per grasp: 1.31
  • Note 1
    See note of Table I for meanings of abbreviations.
  • Note 2
    200+50 – A model trained using the 200 rectangular cuboids dataset plus 50 snack bar dataset; 300+50 – A model trained using the 300 rectangular cuboids dataset plus 50 snack bar dataset;
@@
V-D Experiment 4: Unseen shapes
We also performed experiments to evaluate the generalization ability of the learned policy across different object geometries. Specifically, we tested the trained policy on L-shaped objects, cubes, and hexagons. The L-shaped objects consist of two connected arms, each measuring 60 mm in length. The cross-sectional edge length is 20 mm. The cube and hexagonal objects both have edge lengths of 20 mm. These dimensions were carefully selected to test the robustness of the policy under large variations in object shape and structure.
Table V shows the results without fine-tuning. In this case, the model trained using the 300 dataset was directly applied. Although many grasps were successful, no instances of simultaneous multi-object grasping were observed and none of the trials were fully completed. In all successful grasps, the robot grasped only a single object. The hexagon shape had the lowest performance, likely due to its significant geometric deviation from the training objects. The success rate for the cube was similarly bad, as the robot continuously hesitates around the tube and failure to move forward to grasping, leading to stagnation failures. In contrast, the L-shaped objects achieved the best performance. However, some trials involving L-shapes resulted in notably long grasping durations. These delays primarily occurred when one arm of the L-shape was inside the gripper jaws, causing the robot to hesitate between adjusting toward the other arm or attempting to grasp. This ambiguity led to repeated stalling and prolonged execution.
Table VI shows the results after fine-tuning with an additional 50 demonstrations tailored to the specific object shapes. As expected, the fine-tuned model exhibited improved performance: despite the small amount of additional data, the robot acquired the ability to group objects based on their spatial configuration and perform simultaneous multi-object grasping. These results suggest that even a modest amount of task-specific adaptation can substantially enhance the model’s ability to generalize and execute more complex manipulation behaviors. Fig. 7 shows some exemplary results of the trials. Full videos can be found in the supplementary file.
TABLE V: Unseen objects without fine-tuning
ID L-shape Cube Hexagon
#G Seq. C R #G Seq. C R #G Seq. C R
1 2 (1,1) N P 1 (1) N P 0 - N P
2 1 (1) N P 0 - N P 0 - N P
3 2 (1,1) N P 2 (1,1) N P 0 - N P
4 4 (1,1,1,1) N P 0 - N P 1 (1) N P
5 4 (1,1,1,1) N P 0 - N P 1 (1) N P
  • Note
    #S – Number of successful grasps; #F – Number of failed grasps. All successful grasps involved only a single object. No instances of simultaneous multi-object grasping were observed. The grasping time for each object is shown in the parentheses.
[…]
TABLE VI: Unseen objects with fine-tuned models
ID L-shape Cube Hexagon
#P #G Seq. C R #P #G Seq. C R #P #G Seq. C R
1 0 0 - N P 3 3 (2,2,1) N P 4 4 (2,2,1,1) Y -
2 5 5 (1,1,2,1,1) Y - 3 3 (2,2,2) Y - 1 0 - N G
3 2 1 (1) N G 2 1 (1) N G 4 4 (1,2,2,1) Y -
4 3 2 (2,1) N G 1 1 (2) N P 4 4 (1,2,2,1) Y -
5 4 4 (2,2,1,1) Y - 4 4 (2,1,2,1) Y - 1 1 (2) N P
St 40%, 53%, 1.33 40%, 67%, 1.67 60%, 67%, 1.54
  • Note
    See Table I for meanings of abbreviations and St values.
@@ VI Conclusions and Future Work
In this study, we proposed a diffusion policy-based imitation learning framework for grouping and grasping multiple objects. Experiments show that the robot can flexibly perform pushing, grouping, and grasping across various object configurations, and the pre-trained policy can be fine-tuned for new objects. While effective, the method faces limitations when handling large numbers of objects, scaling to larger workspaces, or adapting to diverse shapes and textures. Future work will focus on exploring the impact of expert strategies on data efficiency, improving data collection, and integrating analytical techniques to enhance robustness and generalization.